OpenStudy (anonymous):

Stephanie is student teaching at her local high school. She has to help her instructor write a test. She has to write a radical equation where the solution is extraneous. She also has to write a radical equation where the solution is non-extraneous. Help Stephanie write one radical equation where the solution is extraneous and another equation where the solution is non-extraneous. Using complete sentences, explain each step when solving to justify your examples.

OpenStudy (anonymous):

I made the non extraneous solution but the extraneous solution i'm getting confused about.

OpenStudy (anonymous):

x^2+4=81

OpenStudy (anonymous):

why negative?

OpenStudy (precal):

sorry my computer is not cooperating

OpenStudy (anonymous):

it's fine and x^2 and 4 are square roots

OpenStudy (precal):

\[\sqrt{7x-1}+8=0\]

OpenStudy (precal):

\[\sqrt{7x-1}+8=0\]

OpenStudy (precal):

ok this one will work, let me explain

OpenStudy (precal):

subtract 8 to both sides

OpenStudy (anonymous):

square both sides

OpenStudy (precal):

\[\sqrt{7x-1}=-8\]

OpenStudy (anonymous):

7x-1=64

OpenStudy (precal):

finish solving it

OpenStudy (anonymous):

x=9

OpenStudy (precal):

no

OpenStudy (anonymous):

7x=65

OpenStudy (precal):

7x-1=64 +1 +1 7x=65 divide 7 to both sides x=65/7

OpenStudy (precal):

ok now sub into original equation

OpenStudy (anonymous):

sry i subtracted instead of added

OpenStudy (precal):

no problem, test it

OpenStudy (anonymous):

\[\pm 8 = 8\]

OpenStudy (anonymous):

-8 i meant

OpenStudy (precal):

no

OpenStudy (precal):

\[\sqrt{7\left( \frac{ 65 }{ 7 } \right)-1}+8=0\]

OpenStudy (precal):

\[\sqrt{65-1}+8=0\] \[\sqrt{64}+8=0\] 8+8=0 16=0 false therefore solution is extraneous

OpenStudy (anonymous):

oh, thx, may i have the original solution

OpenStudy (precal):

x=65/7

OpenStudy (precal):

ok use the one I left

OpenStudy (anonymous):

And here's a medal

OpenStudy (precal):

thanks hope that help you

OpenStudy (anonymous):

You're welcome

OpenStudy (anonymous):

May i ask is This an extraneous solution \[\sqrt{x}=3\]

OpenStudy (precal):

no because the square root of 9 is 3 so x=9 is a solution

OpenStudy (anonymous):

so it's non-extraneous

OpenStudy (precal):

its a solution it is not extraneous its either extraneous or a solution extraneous-extra solution that is in error (it doesn't work)

OpenStudy (anonymous):

Ok i was hoping for it to be non extraneous

OpenStudy (precal):

non is not

OpenStudy (anonymous):

i know